OpenStudy (anonymous):

does anyone know how to do this math problem Solve the linear inequality 7/11x+5/11>-2/11x-2/33 the solution set is =

OpenStudy (ivettef365):

you need to combine like terms as follows: 7 2 2 5 ----- x + ----- x = - ----- - ----- 11 11 33 11 add both fractions

OpenStudy (anonymous):

oh so add 2/33 -5/11

OpenStudy (anonymous):

or no

OpenStudy (anonymous):

add what

OpenStudy (ivettef365):

ok 7/11 + 2/11 = 9/11 so you have and -2/33 + (-5/11) = -17/33 9/11 x = - 17/33

OpenStudy (ivettef365):

Do you have choices?

OpenStudy (anonymous):

oh so the answer is -17/33

OpenStudy (anonymous):

no choices

OpenStudy (ivettef365):

No, that is not the answer now you need to isolate the x

OpenStudy (anonymous):

okay how

OpenStudy (ivettef365):

multiply both sides by 11/9

OpenStudy (anonymous):

okay

OpenStudy (ivettef365):

so you have 11 9 17 11 ---- x --- x = - ---- x ---- 9 11 33 9

OpenStudy (ivettef365):

what did you get ?

OpenStudy (ivettef365):

actually is not (=) is (>)

OpenStudy (anonymous):

huh

OpenStudy (anonymous):

i got 70

OpenStudy (ivettef365):

ok, when you multiply 11/9 x 9/11 that is equal to 99/99 which is equal to 1, so the x isolates so you have x> -17/18

OpenStudy (anonymous):

okay and then you add 17/18

OpenStudy (ivettef365):

no, that is the answer and remember is negative x> - 17/18

OpenStudy (anonymous):

oh okay thank you